A gas sample is initially at 3.3 atm and 8.7 L. If the volume is adjusted to 7.7, what is the new pressure of the gas sample? (Assume the sample is maintained in a closed container with constant temperature).


1
Expert's answer
2021-12-14T06:15:02-0500

P1 = 3.3atm

V1 = 8.7L

V2 = 7.7L

P2 = ?

Using Boyle's law

PV=k(constant)

P1V1 = P2V2

3.3*8.7= P2*7.7

28.71 = P2*7.7

P2= 28.71/7.7

P2 = 3.73atm

The answer is consistent with Boyle's law because when volume decreases, pressure increases(inverse relation)